What is 3 1/3x + 1 5/6 = 4 1/2?
Answer by rfer(16322)

3 1/3x+1 5/6=4 1/2
10/3x+11/6=9/2
10/3x=9/2-11/6
10/3x=27/6-11/6
10/3x=16/6
x=16/6*3/10
x=48/60
x=4/5